OpenStudy (anonymous):

2r - (5 - r) = 13 + 2r. What is r?

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

2r - 5 +r = 13 + 2r => -5 + r =13 => r = 18

OpenStudy (anonymous):

^ There Right Bad Math on my part...